Design a standalone (off grid) PV system for the households in the village so that the following electrical appliances can be utilized: • 1 x 18 W fluorescent lamp running for 4 hours per day • 1 x 60 W fan run for 2 hours per day • 1 x 75 W refrigerator that runs 24 hours per day with compressor that runs 12 hours and off 12 hours The system will be powered by 12 VDC, 110 W PV module.


1
Expert's answer
2021-09-06T00:48:38-0400

Calculating the power consumption demand

"=80*4+60*2+75*12=1340 W\/h"

The total PV energy needed = "1340*1.31742Wh\/day"

The size of the PV panel

The total watt-peak rating needed by the PV

"=\frac{1742}{3.4}=512.352"

The total number of PV needed "= \frac{512.352}{110}=4.65"

Actual number = 5 modulus

Inester sizing

The total watt of all appliance = 80+60+75 =215